OpenStudy (anonymous):

Solve x 2 - 4x - 7 = 0 by completing the square. What are the solutions? a) {2 + squrt of 11, 2 -squrt of 11} b) {-2 + squrt of 11, -2 - squrt of 11} c) {2 +squrt of 7, 2 +squrt of 7}

OpenStudy (anonymous):

http://www.youtube.com/watch?v=gzm-uhj06q8 migth help it comes from expanding (x+a)^2 which is a perfect square (x+a)(x+a)= x^2+2ax+a^2 is a perfect square A B C x 2 - 4x - 7 = 0 is not PS so we must put it in such a form

OpenStudy (anonymous):

Okay thank you the video helped @timo86m

OpenStudy (anonymous):

the process is called completing the square x 2 - 4x - 7 = 0 always get the constant term on the right which is -7 so we add 7 x 2-4x=7 That move enables us to add a constant term to the left which will make it a comple perfect square. ---- we already estaplished that x^2+2ax+a^2 is a perfect square A B C We got x 2 -4x +? we need to find that constant term A B C we know it is a^2 and we know B=-4 B=2a if we plug in -4=2a then solve for a -4/2=a = -2 we add it to both sides since it is algebra and whatever you do to one side you do to the other x 2 -4x+(-2^2)=7+(-2^2) x 2 -4x+4=7+4 x 2 -4x+4=11 since x 2 -4x+4 is a perfect square represent by (x+a)^2 then (x-2)^2=11 x-2= +-sqrt(11) +- is plus or minus x=+-sqrt(11) +2 x=+-3.3166 +2 so x can be 5.31... or -1.31...

OpenStudy (anonymous):

A